OpenStudy (anonymous):

solve algebraically the system: x-y^2=1 (x-3)^2+y^2=4

OpenStudy (anonymous):

Try it by substitution: \[x-y^2=1~~\Rightarrow~~y^2=x-1\] Plug it into the second equation: \[(x-3)^2+(x-1)=4\\ x^2-5x+4=0\\ (x-4)(x-1)=0\\ ~~~~~~~~~~~~\vdots\] Once you find \(x\), plug the solutions into either equation and solve for the corresponding \(y\).